# ColPali: Efficient Document Retrieval with Vision Language Models

### v3

@@ 5 Results
Unstructured Text only
@@ 5.1 Performance (R1)
We iteratively construct ColPali, starting from an off-the-shelf SigLIP model (Table 2).
@@
BiSigLIP: Improving a strong model. SigLIP [^7] https://huggingface.co/google/siglip-so400m-patch14-384 is a strong vision-language bi-encoder model, pretrained on the English split of WebLI (Chen et al., 2023), a corpus of billions of image-text pairs. We find that SigLIP largely outperforms both Jina CLIP and Nomic-vision on document retrieval tasks. Further fine-tuning the textual component of this model on our document-oriented dataset (BiSigLIP) yields clear improvements across the board, particularly on figure retrieval (ArxivQA) and table retrieval tasks (TabFQuAD).
@@
BiPali: Pairing with a language model. In the PaliGemma model architecture, SigLIP-generated patch embeddings are fed to a text language model to obtain LLM contextualized output patch embeddings. [^8] Note that the SigLIP model used in PaliGemma slightly differs in terms of number patches - 1024 patches for PaliGemma’s vision encoder, and 729 for the standalone SigLIP model. We average pool these representations to obtain a single dense vector, effectively creating a PaliGemma bi-encoder model (BiPali). After fine-tuning on the training dataset, we obtain a model that performs slightly worse in English than the tuned BiSigLIP variant. This can be explained by the fact that contrary to SigLIP, the original PaliGemma is not trained on contrastive matching tasks, but rather on next token prediction. Our contrastive fine-tuning phase on 100K images to transform PaliGemma into a bi-encoder is 5 orders of magnitude smaller than SigLIP’s original contrastive training. However, we see notable improvements in French tasks, indicating that BiPali’s LLM (Gemma 2B) helps multilingual text understanding. This is particularly notable as our training dataset does not contain non-English samples.
@@
ColPali: Adding Late Interaction. One benefit of inputting image patch embeddings through a language model is that they are natively mapped to a latent space similar to textual input (query). This enables leveraging the ColBERT strategy to compute interactions between text tokens and image patches, which enables a step-change improvement in performance compared to BiPali. Results in Table 2 show that our ColPali model also largely outperforms the strong baselines based on Unstructured and captioning, as well as all evaluated text-image embedding models. The difference is particularly stark on the more visually complex benchmark tasks, such as InfographicVQA, ArxivQA, and TabFQuAD representing respectively infographics, figures, and tables. However, text-centric documents are also better retrieved by the ColPali models across all evaluated domains and languages, making our approach the overall best-performing document-retrieval model.
@@
Negative Results. For extensiveness, we also train ColSigLIP, a late interaction variant of the BiSigLIP model but obtain abysmal performances. We attribute this to the large gaps w.r.t. SigLIP’s pre-training, in which only a pooled latent representation is used in the contrastive loss, which does not optimize the representations of individual patch and token embeddings. Similarly, we train a BiSigLIPPaliGemma variant, in which we retrieve the image representations from the SigLIP model that has been further updated by PaliGemma fine-tuning, and use the text representations from PaliGemma’s text model. After fine-tuning on our dataset, performance is severely inferior to SigLIPVanilla which simply encodes with SigLIP’s original text and vision components. This indicates a logical misalignment between SigLIP embeddings, and Gemma embeddings after PaliGemma training. We detail these results in Table 5.
@@ 5.2 Latencies & Memory Footprint
Offline Indexing. (R3) Standard retrieval methods using bi-encoders represent each chunk as a single vector embedding, which is easy to store and fast to compute. However, processing a PDF to get the different chunks is the most time-consuming part (layout detection, OCR, chunking), and using captioning to handle multimodal data will only exacerbate this already lengthy process. On the other hand, ColPali directly encodes pages from their image representation. Although the encoder model is larger than standard retrieval encoders, skipping the preprocessing allows large speedups at indexing [^10] Measures a NVIDIA L4 GPU, averaged on 100 pages, with a batch size of 4 pages for ColPali and 8 text chunks for Bi-Encoders. On average, a page is divided into 2.1 chunks. See subsection B.5. (Figure 3).
Memory Footprint. Our method requires storing a vector per image patch. We project each PaliGemma vector to a lower dimensional space (D=128) to maximize efficiency, leading to a memory footprint of $256$ KB per page (subsection B.4). Importantly, the memory footprint of the naive ColBERT indexing strategy can be drastically improved through compression and clustering mechanisms as proposed in the Performance-optimized Late Interaction Driver (Santhanam et al., 2022).
@@ 5.3 Interpretability
By superimposing the late interaction heatmap on top of the original image, we can visualize the most salient image patches with respect to each term of the query, yielding interpretable insights into model focus zones. As epitomized in Figure 1, we observe ColPali exhibits strong OCR capabilities as both the words "hourly" and "hours" present a high similarity score with the query token <_hour>. We also note particular focus on other non-trivial image features such as the x-axis representing hours being salient. Other visualization examples with similar trends of the model transcending pure OCR are shown in Appendix C.
@@ 6 Ablation study
Figure 4: Relative NDCG@5 performance gain w.r.t. the default ColPali (1024 patches). TabFQuAD fine-tuning measures the performance difference on the TabFQuAD task after the introduction of targeted data in the training set. All other results refer to performance deltas averaged on all ViDoRe tasks.
Should we scale models or patch numbers ?
@@
We train a variant of PaliGemma with half the number of image patches (512). While there is a clear performance degradation w.r.t. to the 1024-patch ColPali model (Figure 4), memory usage is much lower. [^11] While another PaliGemma variant exists with 2048 patches, the different training datamix and the large memory requirements make this model impractical for both training and inference time. As an alternative to PaliGemma, we train Idefics2-8B (Laurençon et al., 2024), a VLM with a similar architecture and based on a Mistral-7B (Jiang et al., 2023) language backbone and a SigLIP vision encoder paired with a perceiver resampler. The most notable differences with PaliGemma lie in the size of the language model (2B and 7B resp.) and the number of image patches (between 512 and 2048 for PaliGemma, and 64 post-resampling for Idefics2121212With the option of adding 4 sub-image crops of 64 tokens each to the sequence, for a total of 320 tokens). Our results (Figure 4) suggest language model size has a strong impact on performance, and along with the trained resampler enables more efficient representations for smaller numbers of image embeddings - ColIdefics2 with 64 patches edges out ColPali with 512 patches. Scaling the number of patches of the smaller ColPali model from 512 to 1024, enables largely surpassing the 60-patch ColIdefics2 while being about twice as fast in terms of training and inference latency. These results suggest there are tradeoffs between performance (R1), latencies during online querying (R2) and offline indexation phases (R3), and index memory size.
Should we fine-tune the vision component?
We run our contrastive finetuning on a ColPali model in which we also train the vision encoder and the projection layer. Results in Figure 4 show this leads to no significant improvements.
Do "query augmentation" tokens help?
@@
In ColBERT, special tokens are concatenated to the input query to serve as soft query augmentation buffers. Training without these tokens, we observe no significant performance difference (Figure 4) in the English benchmarks. However, performance on the French tasks seems to improve (Table 5).
Is the Pairwise CE loss best?
@@
Training with an in-batch negative contrastive loss, instead of the pairwise CE loss that only considers the hardest negative sample, leads to a slight performance degradation ($-2.4\%$) on the aggregated benchmark.
Can the model adapt to new tasks?
@@
Contrary to more complex multi-step retrieval pipelines, ColPali can be trained end-to-end, directly optimizing the downstream retrieval task which greatly facilitates fine-tuning to boost performance on specialized domains, multilingual retrieval, or specific visual elements the model struggles with. To demonstrate, we add 1552 samples representing French tables and associated queries to the training set. This represents the only French data in the training set, with all other examples being kept unchanged. We see significant NDCG@5 improvements (Figure 4) and even starker Recall@1 gains ($+6.63\%$) on the TabFQuAD benchmark, with no performance degradation on the rest of the benchmark tasks ($+0.34\%$).
@@ 7 Conclusions
Through the conception of a new benchmark ViDoRe, we established the limits of both modern industrial document retrieval pipelines and off-the-shelf image-text contrastive models for visually rich document retrieval. We introduced ColPali, a novel retrieval model that leverages the latest generative Vision Language models to create highly performing multi-vector embeddings purely from visual document features. ColPali largely outperforms the best existing document retrieval methods while enabling faster corpus indexing time and maintaining low querying latencies, suggesting a very high potential for industrial document retrieval applications. We hope to encourage future work by publicly releasing the ViDoRe benchmark and all models and baselines from our study.
Future Work. Further performance gains could be obtained by exploring sub-image decomposition (Liu et al., 2023a), optimal image patch resampling strategies (Laurençon et al., 2024), or hard-negative mining. Subsequently, our vision is to combine visual retrieval and visually grounded query answering to create RAG systems that purely function from visual features. An interesting line of research could be attempting to generate answers leveraging information stored in the indexed multi-vector patch embeddings.
Limitations
Focus. In this work, we evaluate models on document retrieval tasks, covering several modalities (figures, text, tables, infographics). We however primarily focus on PDF-type documents, and evaluating systems on image retrieval with documents stemming from web page screenshots or hand-written documents might be an interesting generalization. We also focus on high-resource languages (English and French) and although we have shown the capacity of the ColPali model to generalize to languages outside of its fine-tuning set, it is unclear how the model would perform on languages that are not as represented in the model’s language backbone. Finally, our setup assumes relevant documents exist, but abstention methods for Information Retrieval systems might be interesting to explore in more practical settings in which confidence estimation might be important (Gisserot-Boukhlef et al., 2024).
Support. This work relies on multi-vector retrieving derived from the ColBERT late interaction mechanism. Although some vector databases support late interaction engines [^13] Vespa Engine, RAGatouille, QDrant, colbert.ai, many widely used vector retrieval frameworks do not propose native multi-vector support, and some engineering infrastructure efforts may be required to adapt them to work with ColPali (or ColBERT) models.
Data. In the creation of ViDoRe, we partially rely on synthetic query generation based on a commercial large language model, which may induce some amount of bias in the generated queries. To compensate for this, we have iterated on the prompting strategy and given real query examples to the models to help ground generation in realistic settings. We have further manually verified all synthetic queries through a lengthy process to validate their relevance and their quality. Our benchmark also includes many benchmark tasks with no synthetic data, and result trends observed between all tasks are correlated, further confirming the coherence of our benchmark design.
Ethical Considerations
Carbon Footprint. Our work fully leverages prior pretrained models and training is not particularly compute-intensive. Furthermore, we rely on low-rank adapters to further reduce the computational resources needed, both during training and for storage. Overall, a training run represents about 40 hours of Mi250x AMD GPUs. Our experiments, in total, represent 1405 Mi250x GPU hours from highly efficient compute clusters running on low-carbon nuclear energy, representing a total of around 15kg CO2 eq.
Impact. We believe our work could have a strong impact on improving industrial document retrieval systems. Our method is efficient, performs well, and the additional support towards visually rich information from documents could go a long way in unlocking knowledge sources previously difficult to index or query.
Resource Release. For transparency, and to foster future work, we release our comprehensive benchmark under open license and host a public leaderboard [^14] https://huggingface.co/spaces/vidore/vidore-leaderboard. Our models are released under the same usage license as the base model (Gemma Research license for ColPali, Apache2.0 for ColIdefics2) and should be used as intended by the VLM license.
Acknowledgements
This work is partially supported by Illuin Technology, and by a grant from ANRT France. This work was performed using HPC resources from the CINES ADASTRA through Grant 2024-AD011015443. We extend our warm thanks to Jonathan Dong, Caio Corro, Victor Pellegrain and Ender Konukoglu for their valuable feedback on the paper.
@@ A.2 Practical Datasets
Web-Crawler. We implemented a web crawler to efficiently collect large volumes of documents related to a given topic. The crawler is seeded with a user-defined query (e.g. "artificial intelligence") and then uses GPT-3.5 Turbo to brainstorm related topics and subtopics. This query augmentation strategy aims at both broadening and deepening the search. GPT-3.5 Turbo is further used to generate diverse search queries from each subtopic. This query set is then consumed by a pool of parallel workers whose job is to fetch the associated most relevant documents. We use SerpAPI [^15] https://serpapi.com/ along with a filetype filter (PDF documents only) to programmatically scrape Google Search rankings. Each file is hashed and stored in a Bloom filter (Bloom, 1970) shared among workers to avoid duplicate documents in the final corpus. Unique scraped files are downloaded, and inserted into a SQLite database along with additional metadata.
Datamix. Using the web crawler, we collected approximately 1,000 documents for each of the following four seeds: "energy", "government reports", "healthcare industry", and "artificial intelligence". These seeds were meticulously hand-picked to align with real-use cases for retrieval models and visually rich pages. We also removed all documents containing any private information. At this stage, we randomly selected 900 files for the training set and 100 files for the test set, ensuring that data leakage into the test set was avoided during subsequent processing steps.
@@
[…]

### v4

@@ 5 Results
Unstructured text-only
@@ 5.1 Performance (R1)
We show performance is achieved iteratively through the combination of three factors; (1) a carefully crafted task-specific dataset, (2) pairing a pretrained LLM to a vision model to better leverage text semantics from the image, and (3) using multi-vector embeddings rather than a single vector representation to better capture the vast amount of visual information present in a document.
@@
Fine-tuning a Vision Model on a document retrieval oriented dataset: BiSigLIP. SigLIP [^9] https://huggingface.co/google/siglip-so400m-patch14-384 is a strong vision-language bi-encoder producing single vector embeddings, and pretrained on billions of image-text pairs from the English split of WebLI (Chen et al., 2023). Further fine-tuning the textual component of this model on our document-oriented dataset (BiSigLIP) yields clear improvements across the board, particularly on figure retrieval (ArxivQA) and table retrieval tasks (TabFQuAD).
@@
Feeding image patches to a LLM: BiPali. In the PaliGemma model architecture, SigLIP-generated patch embeddings are fed to a text language model and we can obtain LLM contextualized output patch embeddings. [^10] Note that the SigLIP model used in PaliGemma slightly differs in terms of number patches - 1024 patches for PaliGemma’s vision encoder, and 729 for the standalone SigLIP model. This technique aligns the image token representations with the text token embeddings in the LLM’s embeddings space, and augments the vision model embeddings with the language model’s text understanding capabilities. We average pool these representations to obtain a single dense vector, effectively creating a PaliGemma bi-encoder model (BiPali). After fine-tuning on the training dataset, we obtain a model that performs slightly worse in English than the tuned BiSigLIP variant. [^11] This can be explained by the fact that contrary to SigLIP, the original PaliGemma is not trained on contrastive matching tasks, but rather on next token prediction. Our contrastive fine-tuning phase on 119K images to transform PaliGemma into a bi-encoder is 5 orders of magnitude smaller than SigLIP’s original contrastive training. However, we see notable improvements in French tasks, indicating that BiPali’s LLM (Gemma 2B) helps multilingual text understanding. This is particularly notable as our training dataset does not contain non-English samples.
@@
Leveraging Multi-Vector Embeddings through Late Interaction: ColPali. One benefit of inputting image patch embeddings through a language model is that they are natively mapped to a latent space similar to the textual input (query). This enables leveraging the ColBERT strategy to construct one embedding per image patch token, and at inference compute all interactions between text tokens and image patches, resulting in a step-change improvement in performance compared to BiPali. Results in Table 2 show that our ColPali model also largely outperforms the strong baselines based on Unstructured and captioning, as well as all evaluated text-image embedding models. The difference is particularly stark on the more visually complex benchmark tasks, such as InfographicVQA, ArxivQA, and TabFQuAD, respectively representing infographics, figures, and tables. However, text-centric documents are also better retrieved by the ColPali models across all evaluated domains and languages, making our approach the overall best-performing document-retrieval model.
@@
Negative Results. For extensiveness, we also train ColSigLIP, a late interaction variant of the BiSigLIP model but obtain abysmal performances. We attribute this to the large gaps w.r.t. SigLIP’s pre-training, in which only a pooled latent representation is used in the contrastive loss, which does not optimize the representations of individual patch and token embeddings. Similarly, we train a BiSigLIPPaliGemma variant, in which we retrieve the image representations from the SigLIP model that has been further updated by PaliGemma fine-tuning, and use the text representations from PaliGemma’s text model. After fine-tuning on our dataset, performance is severely inferior to SigLIPVanilla which simply encodes with SigLIP’s original text and vision components. This indicates a logical misalignment between SigLIP embeddings, and Gemma embeddings after PaliGemma training. We detail these results in subsection C.1.
@@ 5.2 Latencies & Memory Footprint
Offline Indexing. (R3) Standard retrieval methods using bi-encoders represent each chunk as a single vector embedding, which is easy to store and fast to compute. However, processing a PDF to get the different chunks is the most time-consuming part (layout detection, OCR, chunking), and using captioning to handle multimodal data will only exacerbate this already lengthy process. On the other hand, ColPali directly encodes pages from their image representation. Although the model is larger than standard retrieval encoders, skipping the preprocessing allows large speedups at indexing [^13] Measures a NVIDIA L4 GPU, averaged on 100 pages, with a batch size of 4 pages for ColPali and 8 text chunks for Bi-Encoders. On average, a page is divided into 2.1 chunks. See subsection B.4. (Figure 2). As pages are embedded end-to-end in single forward pass, the VRAM usage depends exclusively on the sequence length (number of patches per image) which is fixed as well, enabling efficient batching strategies to fully leverage hardware acceleration. ColPali also benefits from most LLM efficiency improvements introduced in the ecosystem such as Flash Attention (Dao, 2023).
Storage Footprint. Our method requires storing a vector per image patch, along with 6 extra text tokens “Describe the image” concatenated to image patches. We project each PaliGemma vector to a lower dimensional space ($D=128$) to maximize efficiency, leading to a memory footprint of $257.5$ KB per page (subsection B.3). Importantly, the memory footprint of the naive ColBERT indexing strategy can be drastically improved through compression and clustering mechanisms (Santhanam et al., 2022; Clavié et al., 2024).
Token pooling. Token pooling (Clavié et al., 2024) is a CRUDE-compliant method (document addition/deletion-friendly) that aims at reducing the sequence length of multi-vector embeddings. For ColPali, many image patches share redundant information, e.g. white background patches. By pooling these patches together, we can reduce the amount of embeddings while retaining most of the page’s signal. Retrieval performance with hierarchical mean token pooling on image embeddings is shown in Figure 3 (left). With a pool factor of 3, the total number of vectors is reduced by $66.7\%$ while $97.8\%$ of the original performance is maintained. We note that the Shift dataset—composed of the most text-dense documents—is a clear outlier, showcasing more information dense documents contain less redundant patches and may be prone to worse performance degradation with such pooling techniques.
@@ 5.3 Interpretability
Figure 3: (Left: Token Pooling) Relative performance degradation when reducing the number of stored embeddings per document. (Right: Interpretability) For each term in a user query, ColPali identifies the most relevant document image patches (highlighted zones) and computes a query-to-page matching score.
By superimposing the late interaction heatmap on top of the original image, we can visualize the most salient image patches with respect to each term of the query, yielding interpretable insights into model focus zones. As epitomized in Figure 3 (right), we observe ColPali exhibits strong OCR capabilities as both the words “hourly” and “hours” present a high similarity score with the query token <_hour>. We also note particular focus on other non-trivial image features such as the x-axis representing hours being salient. Other visualization examples are shown in Appendix D.
@@ 6 Ablation study
We run various ablations to better understand the mechanisms at play. By default, result deltas reported below refer to nDCG@5 values averaged over all ViDoRe tasks. Detailed results in C.2.
@@
Tradeoffs between model size and the number of image patches. We train a variant of PaliGemma with half the number of image patches (512). While we observe a clear performance degradation with respects to to the 1024-patch ColPali model ($-24.8$ nDCG@5), memory usage is much lower. As an alternative to PaliGemma, we train Idefics2-8B (Laurençon et al., 2024b), a VLM with a similar architecture and based on a Mistral-7B (Jiang et al., 2023) language backbone and a SigLIP vision encoder paired with a perceiver resampler. The most notable differences with PaliGemma lie in the size of the language model (2B and 7B resp.) and the number of image patches (between 512 and 2048 for PaliGemma, and 64 post-resampling for Idefics2141414With the option of adding 4 sub-image crops of 64 tokens each to the sequence, for a total of 320 tokens). Our results suggest better language models enable more efficient representations of image embeddings - ColIdefics2 with 64 patches largely outperforms out ColPali with 512 patches (+20.1 nDCG@5). However ColIdefics2 (64) remains less accurate than ColPali (1024) ($-4.7$ nDCG@5) while being about twice as slow in terms of training and inference latency. These results suggest there are tradeoffs between performance (R1), latencies during online querying (R2) and offline indexation phases (R3), and index memory size.
Unfreezing the vision component. We train a ColPali variant by also backpropagating through and updating the vision encoder and the projection layer. This leads to a slight performance degradation ($-0.7$ nDCG@5). These conclusions may change with larger scales of training data.
@@
Impact of “query augmentation” tokens. In ColBERT, special tokens are concatenated to the input query to serve as soft query augmentation buffers. Training without these tokens, we observe no significant performance difference in the English benchmarks. However, performance on the French tasks seems to improve ($+9.8$ nDCG@G on Shift, $+6.3$ nDCG@5 on TabFQuAD, Table 7).
@@
Impact of the Pairwise CE loss. Training with an in-batch negative contrastive loss, instead of the pairwise CE loss that only considers the hardest negative sample, leads to a slight performance degradation ($-1.6$ nDCG@5) on the aggregated benchmark.
@@
Adapting models to new tasks. Contrary to more complex multi-step retrieval pipelines, ColPali can be trained end-to-end, directly optimizing the downstream retrieval task which greatly facilitates fine-tuning to boost performance on specialized domains, multilingual retrieval, or specific visual elements the model struggles with. To demonstrate, we add 1552 samples representing French tables and associated queries to the training set. This represents the only French data in the training set, with all other examples being kept unchanged. We see clear nDCG@5 improvements ($+2.6$) and even starker Recall@1 gains ($+5$) on the TabFQuAD benchmark, with no performance degradation on the rest of the benchmark tasks ($+0.4$ nDCG@5 overall).
Better VLMs lead to better visual retrievers. As improved VLMs are released, it is interesting to observe if improved performances on generative tasks translate once these models are adapted for image retrieval tasks through ColPali training strategies. We train the recently released Qwen2-VL 2B (Wang et al., 2024b), a SOTA 2 billion parameter generative VLM, with the same data and training strategy, obtaining ColQwen2-VL. To approximately match ColPali’s memory requirements, we limit the number of image patches to 768, slightly less than ColPali’s 1024 patches. We observe clear performance improvements of $+5.3$ nDCG@5 values over ColPali showcasing clear performance correlations between generative benchmarks performance and retrieving metrics.
Out-of-domain generalization. Some of the datasets in the ViDoRe benchmark have train sets, which we have integrated within the ColPali train set (eg. academic tasks). This is standard in embedding models (Wang et al., 2024a; Lee et al., 2024), and while ColPali also exhibits strong performance on tasks in which this is not the case (French data is never seen by the model during training for instance), it remains interesting to evaluate model performance when training is done on a fully disjoint data distribution. We train a ColPali variant solely using the recent DocMatix dataset (Laurençon et al., 2024a), a large scale, synthetically annotated visual document question answering dataset, which we subsample to obtain a comparably-sized train set. Results on ViDoRe show the performance drop is minor ($-2.2$ nDCG@5), still outperforming the closest baseline method by over 12 points. These results showcase ColPali generalizes well outside of its training distribution, and demonstrate that our results are not unreasonably boosted with respect to baselines (BGE-M3) that cannot be fine-tuned on the same data [^15] To train with data resembling the one BGE-M3 models would see at inference time would require running complex extraction pipelines for the more than 100K documents in the training set, notably relying on external proprietary captioning models which is both too costly and lengthy. This is not needed to train vision-based models.
@@ 7 Conclusions
In this work, we introduced the Visual Document Retrieval Benchmark (ViDoRe), which evaluates document retrieval systems in realistic settings involving visually complex documents. We demonstrated that current retrieval pipelines and contrastive vision-language models struggle to efficiently exploit visual information embedded in documents, leading to suboptimal performance. To address this, we presented ColPali, a novel retrieval method that leverages Vision-Language Models to create high-quality, multi-vector embeddings purely from visual document features. ColPali largely outperforms the best existing document retrieval methods while enabling faster corpus indexing times and maintaining low querying latencies, thus circumventing many pain points of modern document retrieval applications. We hope to drive industrial adoption, and to encourage future work by publicly releasing the ViDoRe benchmark, the data, the codebase, and all models and baselines from our work.
Future Work. Beyond performance improvements that could be obtained through better data, backbone models or training strategies, our vision at term is to combine visual retrieval systems and visually grounded query answering to create end-to-end RAG systems that purely function from image features. This idea is supported by concurrent work (Ma et al., 2024) showcasing the strong promises of VLMs for visual QA, and may eventually become a new industrial standard for document processing. In this line of work, reliability is key, and confidence estimation techniques for Information Retrieval methods could become central to implement abstention mechanisms (Gisserot-Boukhlef et al., 2024), and are particularly interesting given the information rich multi-vector scoring mechanisms of late interaction systems. Expanding benchmarking efforts to cover more languages, modalities, and tasks is also a crucial future research direction (Jiang et al., 2024).
Reproducibility Statement
For transparency, reproducibility and to foster future work, we release our training data, model checkpoints (adapters), entire codebase, and complete evaluation benchmark under MIT licenses as detailed in the main paper. We also host a public ViDoRe leaderboard [^16] https://huggingface.co/spaces/vidore/vidore-leaderboard to foster concurrent work in the space. The supplementary material further details training configurations for our models (also specified in HuggingFace model repositories), and dives into the process we used to generate synthetic data, how latency computations are performed, as well as provides further detailed evaluation results.
@@ A.2 Practical Datasets
Web-Crawler. We implemented a web crawler to efficiently collect large volumes of documents related to a given topic. The crawler is seeded with a user-defined query (e.g. “artificial intelligence”) and then uses GPT-3.5 Turbo to brainstorm related topics and subtopics. This query augmentation strategy aims at both broadening and deepening the search. GPT-3.5 Turbo is further used to generate diverse search queries from each subtopic. This query set is then consumed by a pool of parallel workers whose job is to fetch the associated most relevant documents. We use SerpAPI [^17] https://serpapi.com/ along with a filetype filter (PDF documents only) to programmatically scrape Google Search rankings. Each file is hashed and stored in a Bloom filter (Bloom, 1970) shared among workers to avoid duplicate documents in the final corpus. Unique scraped files are downloaded, and inserted into a SQLite database along with additional metadata.
Datamix. Using the web crawler, we collected approximately 100 documents for each of the following four seeds: “energy”, “government reports”, “healthcare industry”, and “artificial intelligence”. These seeds were meticulously hand-picked to align with real-use cases for retrieval models and visually rich pages. We also removed all documents containing any private information.
@@
You are an assistant specialized in Multimodal RAG tasks. The task is the following: given an image from a pdf page, you will have to generate questions that can be asked by a user to retrieve information from a large documentary corpus. The question should be relevant to the page, and should not be too specific or too general. The question should be about the subject of the page, and the answer needs to be found in the page. Remember that the question is asked by a user to get some information from a large documentary corpus that contains multimodal data. Generate a question that could be asked by a user without knowing the existence and the content of the corpus. Generate as well the answer to the question, which should be found in the page. And the format of the answer should be a list of words answering the question. Generate at most THREE pairs of questions and answers per page in a dictionary with the following format, answer ONLY this dictionary NOTHING ELSE: {     "questions": [         {             "question": "XXXXXX",             "answer": ["YYYYYY"]         },         {             "question": "XXXXXX",             "answer": ["YYYYYY"]         },         {             "question": "XXXXXX",             "answer": ["YYYYYY"]         },     ] } where XXXXXX is the question and [’YYYYYY’] is the corresponding list of answers that could be as long as needed. Note: If there are no questions to ask about the page, return an empty list. Focus on making relevant questions concerning the page. Here is the page:
@@
A.3 Training Dataset
The statistics of the train set are given in the following table. The creation of the train set follows the same methodology as in subsection A.2. We made sure that a PDF document cannot have pages in both the training set and the test set to prevent data leakage and that there are no duplicate documents in each split.
Dataset Split Split Size Language Domain
DocVQA 39,463 English Scanned documents from UCSF Industry
InfoVQA 10,074 English Infographics scrapped from the web
TATDQA 13,251 English High-quality financial reports
arXivQA 10,000 English Scientific Scientific Figures from arXiv
Scrapped PDFs 45,940 English Varied PDFs from 3885 distinct URL domains
TOTAL 118,695 English-only Mixed
Table 3: Details on the different splits in the dataset used to train ColPali.